# A Theory of Learning with Autoregressive Chain of Thought

### v1

@@ D.1 Context and Iterated Sparse Linear Thresholds
We discussed how achieving the desiderata in Section 5 posits an implicit requirement on the universal base class that it has low complexity and high context. Below we define the context formally.
Definition 12 (Context).
We say that a base hypothesis class ${\mathcal{F}}\subseteq\Sigma^{\Sigma^{*}}$ over an alphabet set $\Sigma$ has a context size $d$ if for any function $f\in{\mathcal{F}}$ and ${\bm{x}}\in\Sigma^{*}$, we have $f({\bm{x}})=f({\bm{x}}[-d:])$, that is the output $f$ only depends on the last $d$ coordinates of the input.
Below is a bound on the cardinality of such a class in terms of $\mathrm{VCdim}({\mathcal{F}})$. This bound is a generalization of the cardinality bound of linear thresholds on the boolean hypercube (Lemma 4.1) but for more general hypothesis classes with a bounded context window.
Lemma D.1.
Consider any finite $\Sigma$ and a hypothesis class ${\mathcal{F}}\subseteq\Sigma^{\Sigma^{*}}$ of context length $d\in{\mathbb{N}}_{+}$. For any generation length $T\in{\mathbb{N}}_{+}$,
$\log|{\mathcal{F}}|\leq(d+3)\cdot\textnormal{Ndim}({\mathcal{F}})\cdot\log_{2}|\Sigma|.$
For $\Sigma=\{0,1\}$, this simplifies to $\log|{\mathcal{F}}|\leq(d+3)\cdot\mathrm{VCdim}({\mathcal{F}}).$
Proof.
Since ${\mathcal{F}}$ has a context-window length of $d$, its effective domain is $\Sigma^{\leq d}$. Therefore, the effective number of distinct hypotheses over the domain $\Sigma^{\leq d}$ can be bounded
$|{\mathcal{F}}|\leq\Gamma_{{\mathcal{F}}}(|\Sigma^{\leq d}|)\leq\left(|\Sigma|^{2}|\Sigma^{\leq d}|\right)^{\textnormal{Ndim}({\mathcal{F}})}\leq\,|\Sigma|^{(d+3)\textnormal{Ndim}({\mathcal{F}})}\,.$
When combined with Theorem 3.1, we obtain the sample complexity of $O(d\cdot\mathrm{VCdim}({\mathcal{F}}))$, i.e. still linear in terms of the context size $d$, and thus this attempt of having $T$ independent sample complexity bound is not taking us any closer to our goal of achieving universality.
Sparse Linear Predictors.
Here, we discuss how the class of sparse linear thresholds already satisfies two of the three desiderata in Section 5 simultaneously. However, it remains open whether it satisfies the first desideratum of expressive power. Let
@@
As mentioned, it remains open whether ${\sf TM}(\mathtt{S},\mathtt{T})$ can be expressed by $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ with $d,T={\mathcal{P}oly}(\mathtt{T})$ and $k=O(1)\lll d$ When $k=O(1)$, the computational tractability is easy to see, as we can just do brute-force search over all possible support and implement $\textsc{Cons}_{{\sf CoT}}$ in a polynomially bounded runtime. Below is the lemma sample complexity.
Lemma D.2.
On the domain $\{0,1\}^{d}$, we have $\log|\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}|=O(k^{2}+k\log d)$; however, $\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}$ has a context size $d$.
Proof.
It is easy to see that $\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}$ has a context of size $\Omega(d)$ for any $1\leq k\leq d$. It remains to show the bound on the $\mathrm{VCdim}$ of the end-to-end class. We do so by bounding the effective number of $k$-sparse linear predictors $\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}$ when the domain is the boolean hypercube $\{0,1\}^{d}$ and then taking the cardinality-based bound. There are at most $\binom{d}{k}$ subsets of $[d]$, such that for any $k$-sparse ${\bm{w}}\in\mathbb{R}^{d}$ with $\lVert w\rVert_{0}\leq k$, we have ${\rm supp}({\bm{w}}):=\{i:{\bm{w}}[i]\neq 0\}\subseteq{\mathcal{I}}\in\binom{[d]}{k}$. Define ${\mathcal{F}}_{\scriptscriptstyle{d,{\mathcal{I}}}}=\{f_{{\bm{w}},b}:{\rm supp}({\bm{w}})\subseteq{\mathcal{I}}\}$. Once we decide on a subset ${\mathcal{I}}$ of coordinates that are allowed to be non-zero in a linear predictor with $|{\mathcal{I}}|=k$, the effective domain is $\{0,1\}^{k}$, and thus, ${\mathcal{F}}_{\scriptscriptstyle{d,{\mathcal{I}}}}\equiv{\mathcal{F}}_{\scriptscriptstyle{k,{\rm lin}}}$. Therefore,
$\Gamma_{{\mathcal{F}}_{\scriptscriptstyle{d,{\mathcal{I}}}}}(2^{d})=\Gamma_{{\mathcal{F}}_{\scriptscriptstyle{k,{\rm lin}}}}(2^{k})\leq(e2^{k})^{\mathrm{VCdim}({\mathcal{F}}_{\scriptscriptstyle{k,{\rm lin}}})}\leq(e2^{k})^{(k+1)},$
where we used Sauer’s lemma (Lemma A.1) and the fact that $\mathrm{VCdim}({\mathcal{F}}_{\scriptscriptstyle{k,{\rm lin}}})\leq k+1$. Therefore, after fixing a subset, effectively there are $(e2^{k})^{(k+1)}$ different hypotheses in ${\mathcal{F}}_{\scriptscriptstyle{d,{\mathcal{I}}}}$ on $\{0,1\}^{d}$. Since
$\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}=\bigcup_{{\mathcal{I}}\in\binom{[d]}{k}}{\mathcal{F}}_{\scriptscriptstyle{d,{\mathcal{I}}}},$
the total number of effectively different hypotheses in $\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}$ on $\{0,1\}^{d}$ is bounded by
$|\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}|\leq\binom{d}{k}(e2^{k})^{(k+1)}\implies\log|\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}|=O(k^{2}+k\log d)\,.$
Combining with Theorem 3.1 and using the fact that $\textsc{Cons}_{{\sf CoT}}$ is also $\textsc{Cons}_{{\sf e2e}}$, we obtain that when $k=O(1)$, the class $\mathcal{F}_{\scriptscriptstyle d,k,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf CoT}$-learnable with sample complexity $O(\log d)$ in a polynomial runtime, and it has a context of size $d$. We now prove that indeed our universal base class satisfies all three desiderata.
@@
[…]
@@
where the second inequality follows from the definition of $f_{\tau}$ (Algorithm 1), the next one follows from the discussion above, and the last one follows directly from the way a Turing machine works. Therefore, at the end-to-end step with $t=\mathtt{T}$, we have $f_{\tau}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}({\mathsf{Pre}}({\bm{\omega}}))=(s_{\mathtt{T}},b_{\mathtt{T}},a_{\mathtt{T}})\,.$ Applying post-processing step returns $a_{\mathtt{T}}=g_{\langle\mathtt{S},\mathtt{T},\tau\rangle}({\bm{\omega}})$, which is the desired answer returned to the Turing machine with the transition rule $\tau$. This establishes ${\sf TM}(\mathtt{S},\mathtt{T})\subseteq{\mathsf{Post}}\circ{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}\circ{\mathsf{Pre}}$.
@@
By Theorem 3.1, we have that ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}$ is ${\sf CoT}$-learnable by the rule $\textsc{Cons}_{{\sf CoT}}$ on a sample set $S_{{\sf CoT}}$ of size
@@
Finally, we need to show that $\textsc{Cons}_{{\sf CoT}}$ for the class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ on $S_{{\sf CoT}}$ of size $|S_{{\sf CoT}}|=O(\varepsilon^{\scriptscriptstyle{-1}}\left(\mathtt{S}\log\mathtt{S}+\log\delta^{\scriptscriptstyle{-1}}\right))\,$ containing inputs of length at most $n$, is implementable in time ${\mathcal{P}oly}(n,\mathtt{T},\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. Due to Theorem 3.6, it suffices to show that the base class consistency $\textsc{Cons}_{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}$ is solvable in time polynomial in its input. This is clearly true for the implementation of $\textsc{Cons}_{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}$ we provide, concluding the proof of this theorem.
@@ Appendix E Complementary Results from Section 3
In this section, we provide all our complementary results including lower bounds during the discussion in Section 3. We start with the proof of Theorem 3.3 that establishes that a linear dependence on the generation length $T$ in the sample complexity is necessary for a guarantee based on $\mathrm{VCdim}$ of the base class.
E.1 Proof of Theorem 3.3
The proof will follow by using Proposition 2 once we have a construction of a family of base classes with $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=\Omega(T\cdot\mathrm{VCdim}({\mathcal{F}}))$. Formally, we show the following.
@@
Proof of Theorem 3.3.
Use the negative direction of the Fundamental Theorem of Statistical Learning (Proposition 2), and recall that Definition 1 is simply a supervised learning for ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ class. ∎
@@
$f_{{{\vec{{\bm{b}}}}}}({\bm{x}})=\begin{cases}b_{jD+k},&\text{if }{\bm{x}}=0^{*}\underbrace{{\bm{x}}_{i}}_{n\,\text{bits}}\underbrace{b_{k}b_{D+k}\dots b_{(j-1)D+k}}_{j\,\text{bits}},\text{ for }((i-1)\text{ mod }D\equiv k-1),\text{ and }0\leq j\leq(T-2),\\ b_{i},&\text{if }{\bm{x}}=0^{*}\underbrace{{\bm{x}}_{i}}_{n\,\text{bits}}\underbrace{b_{k}\,b_{D+k}\dots b_{(T-2)D+k}}_{(T-1)\text{ bits}},\text{ for some }((i-1)\text{ mod }D\equiv k-1),\\ 0,&\text{otherwise}.\end{cases}$ (41)
In simple words, the hypothesis $f_{{{\vec{{\bm{b}}}}}}:\Sigma^{*}\to\Sigma$ is defined in such a way that on the point ${\bm{x}}_{i}$ such that $(i-1)\text{ mod }D=(k-1)$ for $k\in[D]$ outputs $b_{k},b_{D+k},b_{2D+k},\dots,b_{(T-2)D+k}$ in the next $(T-1)$ steps of auto-regression before outputting $b_{i}$ at the $T^{\mathrm{th}}$ step.
@@
for $\text{for any }i\in[D],\text{ we have }f_{{{\vec{{\bm{b}}}}}}({\bm{x}}_{i})=b_{k}=b_{i}.$
@@
Because otherwise the functions in ${\mathcal{F}}$ just outputs 00 on that point. Then by pigeon-hole argument, there exists $i_{1},i_{2}\in M$ such that
@@
    $\ell_{1}=\ell_{2}<(T-1)$: In this case, the main intuition is that, any predictor $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$, either agrees with the bit pattern of $\tilde{{\bm{b}}}$ on the respective bits and outputs the same bit $b_{\ell_{1}D+k}=b_{\ell_{2}D+k}$ or otherwise, it outputs 00s on both points. So, effectively, we cannot have different outputs on the two points. Formalizing this, consider the target labels $(\tilde{y}_{1},\tilde{y}_{2})=(0,1)$. Then for any predictor $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$
@@
    $\ell_{2}=\ell_{1}=(T-1)$: In this case, in order to shatter $\{\tilde{{\bm{x}}}_{1},\tilde{{\bm{x}}}_{2}\}$, we must have $i_{1}\neq i_{2}$ because otherwise any $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$ will have the identical outputs on both the points. W.l.o.g. let $i_{1}<i_{2}$. Then we have $i_{1}=(j_{1}-1)\cdot D+k$ for some $11\leq j_{1}\leq T-$ 1. Then we shall show that for the target labels $(\tilde{y}_{1},\tilde{y}_{2})=(\neg\tilde{b}_{(j_{1}-1)\cdot D+k},1)$, for any predictor $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$
@@
We now discuss that even for the end-to-end classes that have time-dependent compositions have roughly the same complexity, i.e. up to $\log T$ factor, as the worst-case lower bound from Theorem 3.3. We show that $\mathrm{VCdim}({\mathcal{F}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}})=O(T\cdot\mathrm{VCdim}({\mathcal{F}})\log T)$, essentially the same as $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=O(T\cdot\mathrm{VCdim}({\mathcal{F}})$) up to a $O(\log T)$ factor.
@@
We now show that $\Omega(\operatorname{Ldim}({\mathcal{F}}))$ samples are necessary in the worst case. Our lower bound is non-trivial in the sense that we show this for a family of classes that always have $\mathrm{VCdim}({\mathcal{F}})=1$. First, note that by the negative direction of the Fundamental Theorem of Statistical Learning (Proposition 2), it suffices to show that the VC dimension of the end-to-end class, $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$, is lower bounded by $\operatorname{Ldim}({\mathcal{F}})$, to establish the lower bound in the sample complexity of end-to-end learning of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$. The following theorem establishes this for a non-trivial family of classes.
@@
$\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=D\text{ for any }T>D\text{ even over domain $\Sigma^{n}$ for }n=\lceil\log D\rceil+1~{}.$
@@
$f_{{{\vec{{\bm{b}}}}}}({\bm{x}})=\begin{cases}b_{j+1},&\text{if }{\bm{x}}=0^{*}\underbrace{{\bm{x}}_{i}}_{n\text{ bits}}\underbrace{b_{1}b_{2}...b_{j}}_{j\text{ bits}}\text{ for some }i\in[D],0\leq j<D\\ b_{i},&\text{if }{\bm{x}}=0^{*}\underbrace{{\bm{x}}_{i}}_{n\text{ bits}}b_{1}b_{2}...b_{D}b_{i}^{*}\text{ for some }i\in[D]\\ 0,&\text{otherwise}\end{cases}$
@@
  • For 00: Choose $f_{{{\vec{{\bm{b}}}}}}$ with $b_{1}=0$. Then $f_{{{\vec{{\bm{b}}}}}}({\bm{x}}_{1})=b_{1}=0$
@@
To show $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})\leq D$, suppose for contradiction that we could shatter $D+1$ points $\{{\bm{y}}_{1},\dots,{\bm{y}}_{D+1}\}$. Consider the behavior of any $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$ on these points: for each point ${\bm{y}}_{j}$, after $T>D$ steps: if ${\bm{y}}_{j}$ starts with some ${\bm{x}}_{i}$, follows the path $b_{1},\dots,b_{D},b_{i}^{*}$ and the output is $b_{i}$, otherwise (if it deviates or doesn’t start with any ${\bm{x}}_{i}$), the output is 00.
@@
E.4 End-to-End Sample Complexity map Collapse

### v2

@@ D.1 Context and Iterated Sparse Linear Thresholds
Here, we discuss how the class of sparse linear thresholds (at least with bounded norm) already satisfies two of the three desiderata in Section˜5 simultaneously. However, it remains open whether it satisfies the first desideratum of the expressive power. Let
@@
As mentioned, it remains open whether ${\sf TM}(\mathtt{S},\mathtt{T})\subseteq\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ with $d,T={\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$ and $k\lll d$ . However, the computational tractability is easy to see. The important thing is that it has the property of “high-context” and “low-sample-complexity”. Indeed, the $\sf e2e$} sample complexity is $O(k^{2}+k\log d)$.
We now prove that indeed our universal base class satisfies all three desiderata.
@@
Let ${\bm{z}}=f_{\tau}^{{\sf CoT}\textnormal{-}{t}}({\bm{x}})$. By induction, we know that for any $1\leq i\leq t$, we have $p_{i-1}=\sum_{j<i}{\bm{z}}[\,j\,].{\mathsf{move}}={\mathsf{pos}}[\,i\,]$ is the head position of the machine just before the time-step $i$ and this is where the symbol ${\bm{z}}[\,i\,].{\mathsf{symb}}$ got written. By the way a Turing machine operates, the current head position after the end of $t$ steps is $p_{t}=p_{t-1}+b_{t}={\mathsf{pos}}[-1]+{\bm{z}}[-1].{\mathsf{move}}={\mathsf{npos}}[|{\bm{z}}|]$ (by the induction hypothesis). So our goal is to retrieve the symbol $r_{t+1}=\text{Tape}[\,p_{t}\,]$ currently present at this location, i.e. the symbol that was written most recently at this position on the tape. This exactly corresponds to finding the largest $j^{*}$ such that ${\mathsf{pos}}[j^{*}]={\mathsf{npos}}[|{\bm{z}}|]$ and retrieve ${\bm{z}}[j^{*}].{\mathsf{symb}}$, and if $j^{*}$ does not exist then at no point in the past, the head of the machine has been at the position so the tape must contain $\mathbin{\square}\;$ . Our $\textnormal{read-tape}({\bm{z}})$ operation is exactly implementing this step (˜6). So we indeed correctly find the symbol $r_{t+1}$ that was read by the Turing machine in order to make the next $(t+1)^{\mathrm{th}}$ move using our read-tape operation. Also, the read-tape operation outputs the state at the final location ${\bm{z}}[-1].{\mathsf{state}}$ which is $s_{t}$ by the induction hypothesis. Therefore, $\textnormal{read-tape}({\bm{z}})=(s_{t},r_{t+1})$ is satisfied. Then we indeed have
@@
where the second inequality follows from the definition of $f_{\tau}$ (Algorithm˜1), the next one follows from the discussion above, and the last one follows directly from the way a Turing machine works. Therefore, at the end-to-end step with $t=\mathtt{T}$, we have $f_{\tau}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}({\mathsf{Pre}}({\bm{\omega}}))=(s_{\mathtt{T}},b_{\mathtt{T}},a_{\mathtt{T}})\,.$ Applying post-processing step returns $a_{\mathtt{T}}=g_{\langle\mathtt{S},\mathtt{T},\tau\rangle}({\bm{\omega}})$, which is the desired answer returned to the Turing machine with the transition rule $\tau$. This establishes ${\sf TM}(\mathtt{S},\mathtt{T})\subseteq{\mathsf{Post}}\circ{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}\circ{\mathsf{Pre}}$.
@@
By Theorem˜3.1, we have that ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}^{{\sf e2e}\textnormal{-}{\mathtt{T}}}$ is $\sf CoT$}-learnable by the rule $\textsc{Cons}_{{\sf CoT}}$ on a sample set $S_{{\sf CoT}}$ of size
@@
Finally, we need to show that $\textsc{Cons}_{{\sf CoT}}$ for the class ${\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}$ on $S_{{\sf CoT}}$ of size $|S_{{\sf CoT}}|=O(\varepsilon^{\scriptscriptstyle{-1}}\left(\mathtt{S}\log\mathtt{S}+\log\delta^{\scriptscriptstyle{-1}}\right))\,$ containing inputs of length at most $n$, is implementable in time ${\mathcal{P}\hskip-1.0ptoly}(n,\mathtt{T},\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$. Due to Theorem˜3.6, it suffices to show that the base class consistency $\textsc{Cons}_{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}$ is solvable in time polynomial in its input. This is clearly true for the implementation of $\textsc{Cons}_{{\mathcal{F}}_{\scriptscriptstyle{\mathsf{TM},\mathtt{S}}}}$ we provide, concluding the proof of this theorem.
@@ Appendix E Complementary Results from Section 3
In this section, we provide all our complementary results including lower bounds during the discussion in Section˜3. We start with the proof of Theorem˜3.3 that establishes that a linear dependence on the generation length $T$ in the sample complexity is necessary for a guarantee based on $\mathrm{VCdim}$ of the base class.
E.1 Proof of Theorem˜3.3
The proof will follow by using Proposition˜2 once we have a construction of a family of base classes with $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=\Omega(T\cdot\mathrm{VCdim}({\mathcal{F}}))$. Formally, we show the following.
@@
Proof of Theorem˜3.3.
Use the negative direction of the Fundamental Theorem of Statistical Learning (Proposition˜2), and recall that Definition˜1 is simply a supervised learning for ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ class. ∎
@@
[…]
In simple words, the hypothesis $f_{{{\vec{{\bm{b}}}}}}:\Sigma^{*}\to\Sigma$ is defined in such a way that on the point ${\bm{x}}_{i}$ such that $(i-1)\text{ mod }D=(k-1)$ for $k\in[D]$ outputs $b_{k},b_{D+k},b_{2D+k},\dots,b_{(T-2)D+k}$ in the next (T1)(T-1) steps of auto-regression before outputting $b_{i}$ at the $T^{\mathrm{th}}$ step.
@@
$\text{for any }i\in[D],\text{ we have }f_{{{\vec{{\bm{b}}}}}}({\bm{x}}_{i})=b_{k}=b_{i}.$
@@
Because otherwise the functions in $\mathcal{F}$} just outputs 0 on that point. Then by pigeon-hole argument, there exists $i_{1},i_{2}\in M$ such that
@@
    $\ell_{1}=\ell_{2}<(T-1)$: In this case, the main intuition is that, any predictor $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$, either agrees with the bit pattern of $\tilde{{\bm{b}}}$ on the respective bits and outputs the same bit $b_{\ell_{1}D+k}=b_{\ell_{2}D+k}$ or otherwise, it outputs 0s on both points. So, effectively, we cannot have different outputs on the two points. Formalizing this, consider the target labels $(\tilde{y}_{1},\tilde{y}_{2})=(0,1)$. Then for any predictor $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$
@@
    $\ell_{2}=\ell_{1}=(T-1)$: In this case, in order to shatter $\{\tilde{{\bm{x}}}_{1},\tilde{{\bm{x}}}_{2}\}$, we must have $i_{1}\neq i_{2}$ because otherwise any $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$ will have the identical outputs on both the points. W.l.o.g. let $i_{1}<i_{2}$. Then we have $i_{1}=(j_{1}-1)\cdot D+k$ for some $1\leq j_{1}\leq T-1$. Then we shall show that for the target labels $(\tilde{y}_{1},\tilde{y}_{2})=(\neg\tilde{b}_{(j_{1}-1)\cdot D+k},1)$, for any predictor $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$
@@
We now discuss that even for the end-to-end classes that have time-dependent compositions have roughly the same complexity, i.e. up to $\log T$ factor, as the worst-case lower bound from Theorem˜3.3. We show that $\mathrm{VCdim}({\mathcal{F}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}})=O(T\cdot\mathrm{VCdim}({\mathcal{F}})\log T)$, essentially the same as $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=O(T\cdot\mathrm{VCdim}({\mathcal{F}})$) up to a $O(\log T)$ factor.
@@
We now show that $\Omega(\operatorname{Ldim}({\mathcal{F}}))$ samples are necessary in the worst case. Our lower bound is non-trivial in the sense that we show this for a family of classes that always have $\mathrm{VCdim}({\mathcal{F}})=1$. First, note that by the negative direction of the Fundamental Theorem of Statistical Learning (Proposition˜2), it suffices to show that the VC dimension of the end-to-end class, $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})$, is lower bounded by $\operatorname{Ldim}({\mathcal{F}})$, to establish the lower bound in the sample complexity of end-to-end learning of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$. The following theorem establishes this for a non-trivial family of classes.
@@
[…]
@@
[…]
@@
  • For 0: Choose $f_{{{\vec{{\bm{b}}}}}}$ with $b_{1}=0$. Then $f_{{{\vec{{\bm{b}}}}}}({\bm{x}}_{1})=b_{1}=0$
@@
To show $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})\leq D$, suppose for contradiction that we could shatter $D+1$ points $\{{\bm{y}}_{1},\dots,{\bm{y}}_{D+1}\}$. Consider the behavior of any $f_{{{\vec{{\bm{b}}}}}}\in{\mathcal{F}}$ on these points: for each point ${\bm{y}}_{j}$, after $T>D$ steps: if ${\bm{y}}_{j}$ starts with some ${\bm{x}}_{i}$, follows the path $b_{1},\dots,b_{D},b_{i}^{*}$ and the output is $b_{i}$, otherwise (if it deviates or doesn’t start with any ${\bm{x}}_{i}$), the output is 0.
@@
E.4 End-to-End Sample Complexity may Collapse